OpenStudy (anonymous):

how can we measure the volume of half fill diesel cylindrical tank in horizontal position

OpenStudy (gunsodo):

\[\frac{1}{2}\pi{r^2}h\]

OpenStudy (anonymous):

the tank is cylindrical half fill and in horizontal placed the picture is attached

OpenStudy (gunsodo):

From that picture, the formula will be \[\frac{1}{2}\pi{r^2}L\]

OpenStudy (anonymous):

I have found this formula but cant find my correct answer

OpenStudy (anonymous):

its not half filled it may be fill 33% or what ever we may just know the height

OpenStudy (gunsodo):

I think that the question says that half fill, so it's meaning half.

OpenStudy (anonymous):

I am sending my numerical data to you from the figure we have L = 2300mm, Diamter = 1650mm, h= 665mm

OpenStudy (gunsodo):

Yes, I agree with you. I've to use your formula. Haha

OpenStudy (anonymous):

okz thx for cooperate